One solution of the equation 2x^3+9x+7-6=0 is x=-2, what are the other solutions?
Need help!

Answer by math_tutor2020(3816)


If we plugged x = -2 into the given equation, then,
2x^3+9x+7-6 = 0
2(-2)^3+9(-2)+7-6 = 0
2(-8)+9(-2)+7-6 = 0
-16-18+7-6 = 0
-33 = 0
We end up with a false statement.
This means x = -2 cannot possibly be a solution of this current equation.

It seems like you might have made a typo. Please double-check the polynomial. Thank you.